# EMFusion: Conditional Diffusion Framework for Trustworthy Frequency Selective EMF Forecasting in Wireless Networks

### v1

@@ I Introduction
The rapid proliferation of new wireless communication technologies, such as transmissions at higher frequencies, massive antennas, and reconfigurable intelligent surfaces, has led to an unprecedented increase in the radio-frequency (RF)-based electromagnetic field (EMF) sources [mootoo2025emforecaster, chiaraviglio2021health]. This technological surge has amplified public and regulatory concerns regarding potential human health effects associated with EMF exposure [chiaraviglio2021health, icnirp2020guidelines]. Consequently, accurate EMF monitoring and forecasting is becoming indispensable for regulatory compliance and for optimizing network deployment and resource management in a proactive manner [mootoo2025emforecaster, stam2018comparison].
To date, existing research studies in the domain focused on univariate forecasting of aggregate EMF exposure collected across a large bandwidth [mootoo2025emforecaster]. However, knowing the wideband EMF exposure is often not enough for active network planning or proactive network resource allocation. Instead a detailed frequency selective EMF exposure breakdown is necessary to identify dominant exposure sources and pin-point the EMF contributions from specific technologies (e.g., 3G, 4G, and 5G deployments) and network operators. Frequency selective EMF forecasting enables regulatory compliance with higher granularity, optimize network deployment as well as spectrum and power allocation policies. Nevertheless, due to the scarcity of frequency selective EMF datasets, forecasting EMF contributions at specific transmission frequencies with multiple network operators and/or diverse wireless technologies remain a fundamental challenge.
Along another note, the uncertainty in EMF patterns, driven by fluctuating network traffic, user mobility, and propagation conditions, means that deterministic point forecasts are insufficient due to the significant errors they produce. Probabilistic forecasting approaches can provide a comprehensive understanding of potential EMF exposure levels, including their likelihood, range, and associated uncertainties [huang2025probabilistic].
Over the past decade, deep learning (DL) solutions have achieved remarkable progress in time-series forecasting (TSF), yet the trained models remain limited in their ability to represent uncertainty and complex temporal dynamics. Conventional architectures such as recurrent neural networks (RNNs) and transformers are primarily designed for deterministic prediction, producing single-point forecasts that often overlook the inherent randomness of real-world systems. As a result, these models can yield overconfident estimates and struggle to generalize under noisy, incomplete, or dynamic conditions. Moreover, their autoregressive nature can lead to error accumulation over extended horizons, reducing long-term reliability. Recently, Generative diffusion models (DMs) enable learning the full probability distribution of future trajectories through an iterative denoising process [huang2025probabilistic, sohl2015deep, ho2020denoising]. This approach enables them to generate diverse and realistic forecasts, impute missing or corrupted data, transmit semantic wireless signals with high fidelity [ref:imputation, 10896580], and flexibly incorporate contextual variables such as seasonal, holiday, or working hour trends [borovykh2017conditional]. By combining probabilistic reasoning with stable and interpretable generation, diffusion-based forecasting represents a significant step forward in capturing the uncertainty and complexity of temporal phenomena.
@@
Network Traffic [nikravesh2016mobile] MLP, MLPWD, SVM Commercial LTE trial traffic UV/MV
[di2023multivariate] VAR, RNN LTE-A VoIP traffic UV/MV
[dalgkitsis2018traffic] LSTM RNN Vodafone 4G throughput UV
[habib2024transformer] Transformer 5G/LTE O-RAN aggregate traffic UV
[fiandrino2024aichronolens] LSTM + XAI 4G load and active-user datasets UV
[isravel2024multivariate] VAR, LSTM ubiquitous healthcare traffic MV
Channel Prediction [sone2020wireless] Holt–Winters, LSTM Real enterprise WLAN traffic UV/MV BS Traffic
[zhang2021deep] LSTM (SCP), RNN Synthetic LEO mMIMO CSI UV
QoS Prediction [hameed2022toward] Temporal Transformer Real IoT smart-building datasets UV/MV
[dinaki2021forecasting] BiLSTM–CNN CDN video streaming QoE dataset MV
[colpitts2023short] LSTM, Seq2Seq Rural fixed-wireless LTE KPIs MV Weather
[zhang2025probabilistic] Diffusion + TCN Real DTN-inspired sensor datasets MV Tempreature
EMF Exposure [kiouvrekis2024comparative] k-NN, XGBoost EMF measurements UV Population Density XAI
[bakcan2022measurement] CNN Wideband RF-EMF UV
[pala2021examining] LSTM, RNN Wideband RF-EMF UV
[nguyen2024deep] Transformer, CNN Wideband RF-EMF MV
[mootoo2025emforecaster] Patching, MLP Wideband RF-EMF UV
@@
To this end, this paper proposes EMFusion, a conditional diffusion model (CDM) for multivariate probabilistic forecasting of EMF exposure at different frequencies. The considered frequency-selective EMF datasets are multivariate since each timestamp contains a set of EMF measurements across multiple frequency bands, wireless technologies, and network operators, forming a multidimensional observation vector. Modeling these jointly, rather than independently, enables capturing interdependencies such as shared temporal patterns, correlated behavior, and coupled exposure dynamics. As such, EMFusion is capable to model the joint distribution of multi-frequency EMF exposure trajectories, capture complex inter-frequency dependencies, and provide built-in uncertainty quantification directly from the generative process, eliminating the need for post-hoc calibration wrappers. To our best knowledge, none of the existing research in the domain of wireless networks have applied conditional DMs for TSF in general and for EMF TSF in particular. The specific contributions of this paper can thus be summarized as follows:
@@ II-1 Statistical Methods for TSF
Initial approaches to EMF exposure forecasting relied on classical time-series models. These included autoregressive models like ARIMA [box2015time], as well as methods such as kernel ridge regression and support-vector regression. While computationally lightweight and interpretable, these techniques often assume stationarity and require significant manual feature engineering. Their linear parametrisation also limits their accuracy when the underlying drivers of EMF exposure interact non-linearly across multiple time scales. A representative approach is found in [de2019time], who proposed a model that decomposes wide-band EMF radiation into three distinct components: the average exposure level, daily and half-daily periodic patterns (seasonal behavior), and the temporal correlation among residuals [de2019time]. Their model effectively captured different behaviors between weekdays and weekends by applying separate autoregressive model for weekdays and Sunday, and the other autoregressive model for Saturday [de2019time]. While effective, such methods are tailored to specific observed periodicities and struggle to generalize without manual recalibration.
@@ II-2 Deep Learning for TSF
[nikravesh2016mobile] compared Multi-Layer Perceptrons (MLP) as well as Support Vector Machines (SVM), to predict traffic in commercial Long-Term Evolution (LTE) networks. Similarly, [dalgkitsis2018traffic] developed a univariate forecasting platform for 4G networks using Long Short-Term Memory (LSTM) and Recurrent Neural Networks (RNNs), demonstrating superiority over statistical baselines like ARIMA. More recent studies have shifted toward multivariate and interpretable models. [di2023multivariate] characterized LTE-Advanced Voice-over-IP metrics using Vector Autoregression (VAR) and supervised DL sliding-window frameworks. To enhance interpretability, [fiandrino2024aichronolens] introduced AICHRONOLENS, an Explainable AI (XAI) framework that correlates temporal structures with LSTM error diagnoses. In the context of Software-Defined Networking, [isravel2024multivariate] utilized LSTM to model jointly evolving traffic flows for healthcare applications. Furthermore, for open radio access networks, [habib2024transformer] proposed a Transformer-based Autoformer architecture to predict aggregate cell traffic. While these methods achieve strong deterministic accuracy, they primarily focus on point forecasting. They generally lack explicit mechanisms for data imputation handling missing values and do not provide intrinsic uncertainty quantification essential for trustworthy network planning.
@@
Channel prediction is critical for mitigating aging in high-mobility scenarios. [sone2020wireless] benchmarked various models, including Holt-Winters and XGBoost, against LSTMs for enterprise Wireless Local Area Network traffic, analyzing spatio-temporal Access Point (AP) clustering. In satellite communications, [zhang2021deep] proposed a Satellite Channel Predictor based on LSTM units for Low Earth Orbit (LEO) massive Multiple-Input Multiple-Output (mMIMO) systems. Their approach outperformed Kalman filters and standard RNNs in multi-step Channel State Information (CSI) forecasting. Similar to traffic forecasting, these works rely on deterministic models and do not account for the confidence level of predictions, which is vital for robust resource allocation.
@@
Forecasting QoS involves modeling complex dependencies. [dinaki2021forecasting] introduced a hybrid bidirectional LSTM (BiLSTM) and Convolutional neural network (CNN) architecture for multivariate QoS forecasting in Content Delivery Networks (CDNs). [hameed2022toward] utilized a temporal Transformer encoder to predict QoS metrics in IoT networks, effectively capturing long-term dependencies that RNNs often miss. [colpitts2023short] investigated Sequence-to-Sequence (Seq2Seq) models for rural LTE Key Performance Indicators (KPIs), noting that exogenous features provided minimal gains due to short temporal dependencies.
@@
Research in EMF forecasting has progressed from spatial regression to time-series analysis. [kiouvrekis2024comparative] focused on comparative regression using ensemble methods like Random Forest to rank urban factors affecting EMF levels. For temporal forecasting, [pala2021examining] demonstrated that LSTMs outperform classical statistical models (e.g., ARIMA) in long-term (over a 60-month period) for univariate EMF prediction. Recently, [nguyen2024deep] demonstrated the effectiveness of 1-D CNNs in handling single-step and multi-step EMF forecasting tasks by leveraging translation invariance to identify periodicities in the data. Other research works also adapted CNN designs, utilizing architectures like Radial Basis Function Networks and Generalized Regression Neural Networks to model complex RF-EMF signals from emerging technologies like 5G [al2023machine]. Despite the aforementioned advancements, significant research gaps remain in the domain.
@@
    Approaches like [mootoo2025emforecaster] rely on post-hoc calibration rather than learning the joint probability distribution directly. This decoupling can limit the expressiveness of the uncertainty representation. To address this, recent work has turned to distribution-free methods like conformal prediction. A notable example is EMForecaster [mootoo2025emforecaster], which integrates a powerful PatchTST backbone with a post-hoc conformal calibration layer. This two-stage pipeline delivers statistically valid prediction intervals for EMF levels. However, by decoupling uncertainty quantification from the primary sequence modeling, this approach is not directly applicable to multivariate frequency selective forecasting that adhere to the physical constraints and interdependencies across different frequency carriers, wireless technologies, and operators. Furthermore, EMForecaster [mootoo2025emforecaster] is designed for univariate forecasting, thus a large number of parallel models will be required to handle frequency selective forecasting.
@@
    Current works typically assume clean historical data, failing to address the frequent sensor outages and data gaps inherent in real-world EMF monitoring [nguyen2024deep].
@@ II-3 Generative AI for TSF
In recent years, generative models have emerged as a powerful paradigm within DL for TSF. Representative architectures include Variational Autoencoders (VAEs) [li2022generative], Generative Adversarial Networks (GANs) [goodfellow2014generative, koochali2021if], and, more recently, DMs  [song2019generative, ho2020denoising], which have gained remarkable attention for their ability to capture data uncertainty and generate high-fidelity forecasts. DMs, as latent-variable generative frameworks, learn complex data distributions via a dual process of progressive noise injection (the forward diffusion) and iterative denoising (the reverse process). More recently, DMs have been successfully adapted to time series data [rasul2021autoregressive, tashiro2021csdi], offering a flexible generative framework for learning intricate temporal dependencies and stochastic dynamics. While DiffTCN introduces probabilistic modeling, the majority of QoS research remains deterministic.
In the context of TSF, DMs are typically formulated as conditional generative models, where the goal is to produce future trajectories conditioned on past observations and optional contextual covariates [li2022generative, rasul2021autoregressive, alcaraz2022diffusion]. During training, Gaussian noise is incrementally added to the ground-truth future sequence through the forward diffusion process, while a neural network, conditioned on historical context, is optimized to learn the inverse mapping that progressively denoises the signal. At inference, forecasts are generated by iteratively sampling and denoising noise vectors, effectively drawing from the learned conditional distribution.
@@
    DMs yield fully probabilistic forecasts, enabling uncertainty quantification through multiple stochastic samples and predicting extreme cases in terms of different EMF traffic flows, whereas deterministic models are disadvantaged in capturing uncertainty [yan2025semantic].
@@
    They offer a highly modular design that allows flexible conditioning on exogenous factors such as seasonal indicators, control variables, or multimodal inputs [kollovieh2023predict, yuan2024diffusion, liu2024retrieval].
Empirically, DMs demonstrate competitive or even state-of-the-art results across numerous TSF benchmarks [shen2024multi, liu2024retrieval, li2025diffusion, meijer2024rise, su2025diffusion], positioning them as a cornerstone of next-generation generative forecasting frameworks.
@@ III Multivariate EMFusion: Conditional DM for Frequency Selective EMF Forecasting
This section details our methodology for multivariate EMF TSF leveraging a conditional DM. Given the sequence of historical EMF measurements and corresponding contextual factors, the goal is to model the conditional distribution of the future EMF values. The proposed model is then enhanced with imputation-based resampling to deal with real-world EMF logs where EMF measurements may be irregularly sampled or there are gaps or sensor outages.

### v2

@@ I Introduction
The rapid proliferation of new wireless communication technologies, such as transmissions at higher frequencies, massive antennas, and reconfigurable intelligent surfaces (RISs), has led to an unprecedented increase in the radio-frequency (RF)-based electromagnetic field (EMF) sources [1, 2]. This technological surge has amplified public concerns regarding potential human health effects associated with EMF exposure [2, 3, 4]. Consequently, accurate EMF monitoring and forecasting are becoming indispensable for regulatory compliance, network planning, and resource management [5, 6].
Existing studies primarily rely on stochastic modeling and optimization frameworks to manage exposure. Representative approaches include resource allocation and energy-efficiency maximization under EMF limits [7, 8], the integration of RIS to reduce exposure while maintaining service quality [9], and EMF-aware network planning through optimized base station (BS) placement and configuration [10, 11]. Operational strategies, such as exposure-conscious user association [12], beamforming optimization [13], and adaptive cross-layer protocols [14] further aim at exposure mitigation.
Beyond modeling-based strategies, some recent works have strengthened EMF measurement methodologies and long-term monitoring in the context of 5G deployment. For instance, [15] and [16] developed rigorous procedures for estimating the instantaneous maximum received power from 5G sources and refining exposure assessment at millimeter-wave frequencies. The IEC 62232:2025 standard [17] reports case studies evaluating EMF exposure across typical BS scenarios from 100 kHz to 300 GHz. Longitudinal monitoring data in [18] reveal a gradual increase in exposure at 3.6 GHz over two years in Greece, linked to expanding 5G infrastructure.
Despite these advances in EMF modeling and monitoring, EMF forecasting efforts remain limited. The handful of EMF forecasting studies remain primarily focused on univariate prediction of aggregate (wideband) exposure across large bandwidths [5, 19, 20, 21, 22]. However, wideband exposure alone is often insufficient for regulatory oversight, active network planning, or proactive resource allocation. In practice, a frequency-selective breakdown is required to identify dominant exposure sources and quantify the contributions of specific technologies (e.g., 3G, 4G, and 5G deployments) and network operators. From a regulatory perspective, this granularity is essential. Environmental protection agencies must enforce maximum exposure limits at each frequency band across the territory to safeguard public health [^1] Simulation-based assessments submitted during site authorization are typically conservative and may differ from actual measured exposure levels, as operators often declare maximum radiated power, antenna tilting, or beamforming configurations that exceed those used in operation [4]. Extending monitoring capabilities to frequency-selective forecasting would enable forward-looking compliance verification and transform periodic reporting from retrospective documentation into proactive decision support. Moreover, per-frequency forecasts could establish a feedback loop between regulatory oversight and network management, allowing operators to adjust spectrum allocation, transmission power, or deployment strategies, particularly in sensitive areas, before exposure limits are approached.
Although deep learning (DL) has achieved significant progress in time-series forecasting, conventional architectures such as recurrent neural networks (RNNs) and transformers are mainly designed for deterministic prediction. By producing single-point estimates, they often overlook intrinsic randomness, leading to overconfident forecasts and reduced robustness under noisy or dynamic conditions. Their autoregressive structure may also cause error accumulation over longer horizons. Recently, generative diffusion models (DMs) have emerged as a powerful probabilistic alternative, learning the full distribution of future trajectories through an iterative denoising process [23, 24, 25]. Beyond generating diverse and realistic forecasts, DMs support robust data imputation [26, 27] and flexible integration of contextual variables [28], making them well suited to capture the uncertainty and complexity of frequency-selective EMF dynamics.
@@
Network Traffic [29] MLP, MLPWD, SVM Commercial LTE trial traffic UV/MV
[30] VAR, RNN LTE-A VoIP traffic UV/MV
[31] LSTM RNN Vodafone 4G throughput UV
[32] Transformer 5G/LTE O-RAN aggregate traffic UV
[33] LSTM + XAI 4G load and active-user datasets UV
[34] VAR, LSTM ubiquitous healthcare traffic MV
Channel Prediction [35] Holt–Winters, LSTM Real enterprise WLAN traffic UV/MV BS Traffic
[36] LSTM (SCP), RNN Synthetic LEO mMIMO CSI UV
QoS Prediction [37] Temporal Transformer Real IoT smart-building datasets UV/MV
[38] BiLSTM–CNN CDN video streaming QoE dataset MV
[39] LSTM, Seq2Seq Rural fixed-wireless LTE KPIs MV Weather
[40] Diffusion + TCN Real DTN-inspired sensor datasets MV Tempreature
EMF Exposure [41] LSTM, CNN EMF measurements UV Population Density
[19] k-NN, XGBoost EMF measurements UV Population Density XAI
[42] CNN Wideband RF-EMF UV
[20] LSTM, RNN Wideband RF-EMF UV
[21] Transformer, CNN Wideband RF-EMF MV
[5] Patching, MLP Wideband RF-EMF UV
@@
Building on the need for frequency-selective and probabilistic EMF forecasting under uncertainty, this paper proposes EMFusion, a conditional diffusion model (CDM) for multivariate probabilistic forecasting of EMF exposure across multiple frequencies. The considered frequency-selective EMF datasets are inherently multivariate, as each timestamp comprises measurements across several frequency bands, wireless technologies, and network operators, forming a multidimensional observation vector. Modeling these jointly enables capturing shared temporal patterns, correlated behaviors, and coupled exposure dynamics. Accordingly, EMFusion models the joint distribution of multi-frequency EMF trajectories and provides built-in uncertainty quantification directly through its generative process, eliminating the need for post-hoc calibration.
To the best of our knowledge, conditional diffusion models have not been applied to time-series forecasting in wireless networks in general, nor to EMF forecasting in particular. The specific contributions of this paper can thus be summarized as:
@@ II-1 Statistical Methods for TSF
Initial approaches to EMF exposure forecasting relied on classical time-series models. These included autoregressive models like ARIMA [43], as well as methods such as kernel ridge regression and support-vector regression. While computationally lightweight and interpretable, these techniques often assume stationarity and require significant manual feature engineering. Their linear parametrisation also limits their accuracy when the underlying drivers of EMF exposure interact non-linearly across multiple time scales. A representative approach is found in [44], who proposed a model that decomposes wide-band EMF radiation into three distinct components: the average exposure level, daily and half-daily periodic patterns (seasonal behavior), and the temporal correlation among residuals [44]. Their model effectively captured different behaviors between weekdays and weekends by applying separate autoregressive model for weekdays and Sunday, and the other autoregressive model for Saturday [44]. While effective, such methods are tailored to specific observed periodicities and struggle to generalize without manual recalibration.
@@ II-2 Deep Learning for TSF
[29] compared Multi-Layer Perceptrons (MLP) as well as Support Vector Machines (SVM), to predict traffic in commercial Long-Term Evolution (LTE) networks. Similarly, [31] developed a univariate forecasting platform for 4G networks using LSTM and Recurrent Neural Networks (RNNs), demonstrating superiority over statistical baselines like ARIMA. More recent studies have shifted toward multivariate and interpretable models. [30] characterized LTE-Advanced Voice-over-IP metrics using Vector Autoregression (VAR) and supervised DL sliding-window frameworks. To enhance interpretability, [33] introduced AICHRONOLENS, an Explainable AI (XAI) framework that correlates temporal structures with LSTM error diagnoses. In the context of Software-Defined Networking, [34] utilized LSTM to model jointly evolving traffic flows for healthcare applications. Furthermore, for open radio access networks, [32] proposed a Transformer-based Autoformer architecture to predict aggregate cell traffic. While these methods achieve strong deterministic accuracy, they primarily focus on point forecasting. They generally lack explicit mechanisms for data imputation handling missing values and do not provide intrinsic uncertainty quantification essential for trustworthy network planning.
@@
Channel prediction is critical for mitigating aging in high-mobility scenarios. [35] benchmarked various models, including Holt-Winters and XGBoost, against LSTMs for enterprise Wireless Local Area Network traffic, analyzing spatio-temporal Access Point (AP) clustering. In satellite communications, [36] proposed a Satellite Channel Predictor based on LSTM units for Low Earth Orbit (LEO) massive Multiple-Input Multiple-Output (mMIMO) systems. Their approach outperformed Kalman filters and standard RNNs in multi-step Channel State Information (CSI) forecasting. Similar to traffic forecasting, these works rely on deterministic models and do not account for the confidence level of predictions, which is vital for robust resource allocation.
@@
Forecasting QoS involves modeling complex dependencies. [38] introduced a hybrid bidirectional LSTM (BiLSTM) and Convolutional neural network (CNN) architecture for multivariate QoS forecasting in Content Delivery Networks (CDNs). [37] utilized a temporal Transformer encoder to predict QoS metrics in IoT networks, effectively capturing long-term dependencies that RNNs often miss. [39] investigated Sequence-to-Sequence (Seq2Seq) models for rural LTE Key Performance Indicators (KPIs), noting that exogenous features provided minimal gains due to short temporal dependencies.
@@
Research in EMF forecasting has progressed from spatial regression to time-series analysis. [19] focused on comparative regression using ensemble methods like Random Forest to rank urban factors affecting EMF levels. For temporal forecasting, [20] demonstrated that LSTMs outperform classical statistical models (e.g., ARIMA) in long-term (over a 60-month period) for univariate EMF prediction. Recently, [21] demonstrated the effectiveness of 1-D CNNs in handling single-step and multi-step EMF forecasting tasks by leveraging translation invariance to identify periodicities in the data. Other research works also adapted CNN designs, utilizing architectures like Radial Basis Function Networks and Generalized Regression Neural Networks to model complex RF-EMF signals from emerging technologies like 5G [22]. Despite the aforementioned advancements, significant research gaps remain in the domain.
@@
    Approaches like [5] rely on post-hoc calibration rather than learning the joint probability distribution directly. This decoupling can limit the expressiveness of the uncertainty representation. To address this, recent work has turned to distribution-free methods like conformal prediction. A notable example is EMForecaster [5], which integrates a powerful PatchTST backbone with a post-hoc conformal calibration layer. This two-stage pipeline delivers statistically valid prediction intervals for EMF levels. However, by decoupling uncertainty quantification from the primary sequence modeling, this approach is not directly applicable to multivariate frequency selective forecasting that adhere to the physical constraints and interdependencies across different frequency carriers, wireless technologies, and operators. Furthermore, EMForecaster [5] is designed for univariate forecasting, thus a large number of parallel models will be required to handle frequency selective forecasting.
@@
    Current works typically assume clean historical data, failing to address the frequent sensor outages and data gaps inherent in real-world EMF monitoring [21].
@@ II-3 Generative AI for TSF
In recent years, generative models have emerged as a powerful paradigm within DL for TSF. Representative architectures include Variational Autoencoders (VAEs) [45], Generative Adversarial Networks (GANs) [46, 47], and, more recently, DMs  [48, 25], which have gained remarkable attention for their ability to capture data uncertainty and generate high-fidelity forecasts. DMs, as latent-variable generative frameworks, learn complex data distributions via a dual process of progressive noise injection (the forward diffusion) and iterative denoising (the reverse process). More recently, DMs have been successfully adapted to time series data [49, 50], offering a flexible generative framework for learning intricate temporal dependencies and stochastic dynamics. While DiffTCN introduces probabilistic modeling, the majority of QoS research remains deterministic.
In the context of TSF, DMs are typically formulated as conditional generative models, where the goal is to produce future trajectories conditioned on past observations and optional contextual covariates [45, 49, 51]. During training, Gaussian noise is incrementally added to the ground-truth future sequence through the forward diffusion process, while a neural network, conditioned on historical context, is optimized to learn the inverse mapping that progressively denoises the signal. At inference, forecasts are generated by iteratively sampling and denoising noise vectors, effectively drawing from the learned conditional distribution.
@@
    DMs yield fully probabilistic forecasts, enabling uncertainty quantification through multiple stochastic samples and predicting extreme cases in terms of different EMF traffic flows, whereas deterministic models are disadvantaged in capturing uncertainty [52].
@@
    They offer a highly modular design that allows flexible conditioning on exogenous factors such as seasonal indicators, control variables, or multimodal inputs [53, 54, 55].
Empirically, DMs demonstrate competitive or even state-of-the-art results across numerous TSF benchmarks [56, 55, 57, 58, 59], positioning them as a cornerstone of next-generation generative forecasting frameworks.
@@ III Multivariate EMFusion: Conditional DM for Frequency Selective EMF Forecasting
This section details our methodology for multivariate EMF TSF leveraging a conditional DM. Given the sequence of historical EMF measurements and corresponding contextual factors, the goal is to model the conditional distribution of the future EMF values. The proposed model is then enhanced with imputation-based resampling to deal with real-world EMF logs where EMF measurements may be irregularly sampled or there are gaps or sensor outages. Before detailing the formal mathematical framework, we provide a conceptual overview of the three core pillars of the EMFusion architecture:
  • Diffusion Modeling as Iterative Refinement: Unlike standard models that predict a single value, diffusion modeling treats forecasting as a process of systematic denoising. The model learns to reconstruct a structured, high-fidelity EMF signal from Gaussian noise through multiple refinement steps [25, 23].
  • Cross-Attention as Dynamic Weighting: To ensure forecasts are context-aware, we utilize a cross-attention mechanism. This acts as a dynamic lens, allowing the model to selectively focus on specific historical patterns or external conditions (e.g., working hours or seasons) that are most relevant to the EMF time-series [23, 27].
  • Inpainting as Pattern Completion: We frame the forecasting task as a structural inpainting problem. Much like filling in a missing piece of a photograph, the model treats future EMF values as the missing portion of a temporal image and utilizes the known ”texture” of the past to complete the sequence realistically [26].